OpenStudy (anonymous):

a triangle with side a=? b=12 c=14 Find x ?

OpenStudy (anonymous):

IF YOU ARE GOOD AT ANALYZING POEMS GO TO http://www.twiddla.com/548645

OpenStudy (anonymous):

Pythagorean theorem! \[a^2+b^2=c^2\]\[a^2+144=196\]\[a^2=52\]\[a=\sqrt{52}\]